# Privacy Ripple Effects from Adding or Removing Personal Information in Language Model Training

### v1

@@ 1 Introduction
    We observe the phenomenon of assisted memorization: PII may not be memorized immediately after it is seen, but may be memorized later in training (§5 and Figure 1). We find this is largely influence by $n$-gram statistics.
@@ 2 Related Work
studies when a text is trained on and generated by a model. This is widely studied (Carlini et al., 2019, 2021, 2022b; Lee et al., 2022; Zhang et al., 2023; Ippolito et al., 2023; Biderman et al., 2023a, b; Kudugunta et al., 2024; Nasr et al., 2023; Chang et al., 2023; Ozdayi et al., 2023; Schwarzschild et al., 2024; Duarte et al., 2024; Wang et al., 2024). These works are often focused on the broad phenomenon, and not the nature of the data, e.g., if it were sensitive as in our work. Relatively fewer works have considered this setting. Huang et al. (2022) study if information about specific entites can be extracted; Panda et al. (2024) study if LLM’s can be poisoned to memorize specific PII; Lukas et al. (2023) formalize PII extraction, proposing several attacks and studying the efficacy of various existing defenses; and Lehman et al. (2021) found that extracting sensitive data, using simple techniques, from BERT trained on clinical notes was largely unsuccessful. This line of work has become important for practical privacy and memorization audits (Anil et al., 2023; Gemini Team et al., 2023; Dubey, 2024), which also often include PII memorization evaluations (Gemini Team et al., 2023, 2024; Gemma Team et al., 2024a, b; CodeGemma Team et al., 2024).
@@ 3 Experimental Setup
We fine-tune these models on two datasets. First, we use a modified version of the WikiText-2 dataset Merity et al. (2016) to include unique emails from the Enron dataset [^7] https://www.cs.cmu.edu/enron/. We take the entire WikiText-2 dataset and insert $E$ unique email addresses (herein, emails) into each passage. We concatenate all passages during training and divide them into blocks of 128 tokens. Second, we use the Pile of Law dataset (Henderson et al., 2022). We ensure no emails were already memorized by querying the base models with the same prompts. Lee et al. (2022) found data duplication strongly increases memorization. In our study, all emails occur in the training corpus exactly once.
@@
We closely follow the methodology of Carlini et al. (2021); Nasr et al. (2023). We focus on “extractable memorization” and use ten-token sequences sampled uniformly at random from Common Crawl. We randomly sample a unique set of $25,000$ different prompts for each experiment. We obtain a $256$ token output from the model for each prompt and evaluate it for successful extraction. Our method may lead to false negatives; however, this would only underestimate the PII regurgitation, and, we further believe our diverse and large prompt dataset reasonably captures the regurgitation rates. To further minimize false-negatives, where denoted we also evaluate “discoverable” memorization, where we prompt with the exact prefix the model trained on. We use greedy decoding, or top-$k=40$ sampling when specified.
@@
We primarily use the definition of extractable memorization (and, where denoted, discovered memorization) from  Nasr et al. (2023). Herein, we will refer to a success as an extraction, which is whenever an email is contained both in the training dataset and a language model’s generation. Formally, let $\mathcal{D}$ be the training dataset for a language model $M$. Let $f$ be a chosen sampling scheme that takes an input text prompt $p$ and returns the conditional generation $s=f_{M}(p)$. An email $e^{i}$ is said to be extracted if $e^{i}\in\mathcal{D}$ and $\exists p:e^{i}\in f_{M}(p)$.
@@ 4.2 Experimental Results
We also find that models are constantly forgetting. Throughout the entirety of training (including the beginning and end), many models (see Appendix B for more results on other models and datasets) exhibit a cycle of forgetting and immediate memorization. This result sheds new light into the dynamic view of memorization: which samples are memorized by a model may be more a function of stochasticity than previously thought. The choice of which model to release may play a larger role in determining which samples are memorized, due to which samples were forgotten or re-memorized than previously thought due to the stochasticity in data sampling.
@@
When using our taxonomy to analyze memorizing, we observe that a significant fraction of memorization samples are not classified by these three categories. This leads to an another interesting finding: a lot of memorization is not immediately memorized. In other words, at a given step, other text that was not trained on at this step is now extractable by the model.
@@
Our results in Figure 3 show that LLMs do forget some of the previously memorized PII as training progresses. Prior work has shown that some examples memorized early in training may be forgotten after additional training (Jagielski et al., 2022). Further, we also observe that some forgotten emails get re-extracted when there is $n$-gram overlap between tokens from the email and tokens in the data during further training. This phenomenon is illustrated in Figure 4, which shows how previously extracted samples that the model later forgets can reappear at subsequent checkpoints. Each cell indicates the percentage of emails extracted both by the corresponding checkpoint and the reference checkpoint (diagonal cell). Since each diagonal cell serves as its own reference, its value is always 1.
Figure 3: Tracking memorization throughout training with our taxonomy. The stacked bars show how many newly memorized emails are immediate, retained, and assisted, while red denotes forgotten emails since the last checkpoint. We see large amounts of assisted memorization occurring later in training, underscoring that PII is not always memorized immediately. Takeaway: memorization is more dynamic and stochastic than often assumed, with ongoing cycles of forgotten and newly assisted emails.
Figure 4: Forgotten PII is re-extracted later. The diagonal values $d_{ii}$ represent the total extraction at each checkpoint; off-diagonal cells show which fraction of emails remain memorized at both checkpoints. Takeaway: memorized PII can sometimes slip out of memory, only to reappear once certain overlapping tokens occur in future training steps.
@@ 5 Assisted Memorization: Training on One’s PII Can Reveal Another’s
In Figure 3, we see that a large fraction of memorization is assisted. This is especially true later in training, where we observe that more memorization is assisted than immediate, specifically a mean rate of $0.03$ for assisted compared to $0.01$ for immediate. This finding is not model- or data-specific, as our results in Appendix B (e.g., Figures 16 and 17) show similar trends.
The existence of assisted memorization brings to light a deeper privacy concern. One may expect that data seen earlier is less vulnerable to privacy risks through a form of “recency bias” (implied by forgetting effects). Our findings of assisted memorization, however, show that this may not always be the case; the existence of this effect with sensitive data like PII is of particular concern because it shows that downstream training stages must be careful how they may elicit the extraction of earlier training data. The most common practical scenario for this is in the pre-training/fine-tuning setup that current LLMs undergo. Our results show that fine-tuning even on natural (non-adversarially) constructed training datasets can uncover the extraction of PII in pre-training data. Prior work (Nasr et al., 2023) only showed this may be possible with adversarial constructions. Pragmatically, our results also show that privacy and memorization audits, especially when PII is of concern, should encompass all data in the training history, and not just data from the most recent training stage.
@@ 5.1 Assisted Memorization Is Not Simply Delayed
We know when data samples were first seen from data sampling. Then, we must identify exactly when each email became extractable, as any training beyond this may lead to forgetting. Given that we only checkpoint our models every 10% of training, for efficiency, we do not have this a priori. To determine this, we use a binary search performing an extraction test on each iteration of the search. This significantly reduces the overhead as the extraction test is expensive (recall we prompt the model thousands of times as described in §3).
@@ 5.2 Assisted Memorization Is Triggered by Training on Specific $n$-grams
Next, we ask: when multiple emails share a firstname, why might a particular email with a different lastname get assisted memorized over another? For example, why might john.mccarthy@gmail.com be memorized over john.williams@gmail.com. We train a simple logistic regression model on features capturing $n$-grams overlaps, last-name counts, and domain counts for all assisted memorized emails (positives) and those not memorized (negatives). More details are in Appendix C.
Our logistic regression model is trained to predict assisted memorized emails from a dataset consisting of these emails labeled as positive, and other emails sharing the same firstname but a different lastname as negatives. We use a standard 5-way cross validation setup with 10 trials. Full details are in Appendix C. The model achieves a precision of 0.937 and recall of 0.874 indicating high success.
In Figure 5, we visualized the logistic regression model’s score against the email likelihood from $M$, computed against the successful prompt that led to extraction. This shows that assisted memorization emails tend to be well classified from these simple features. We observe that $n$-gram statistics were the most important feature, further supporting our conclusions above (see Table 1 of Appendix C where we report the feature weights).
@@ 6.1 Contributing More Data via Opt-ins
Next, we train ten distinct models $M_{1}$ to $M_{10}$, where $M_{i}$ is trained on $D_{10i\%}$ for three epochs, following the same training setup described in Section 3. We highlight that the only change between these models is the additional emails. Otherwise, we use the same training process and the same prompts for all models when decoding.
@@
We report the results of our experiment in Figure 6, for models finetuned for three epochs (more results in Appendix D). We highlight two major findings.
First, we find that the number of extracted emails increases substantially with the amount of PII contained in the model’s finetuning set. This can be seen on the diagonals of Figure 6, which show the total amount of PII extracted from the relevant model. For top-$k$ sampling, we see that 283 emails are extracted from $M_{10}$, compared to only 57 at $M_{5}$, which was trained on half as many emails—the increase in extraction from top-$k$ sampling is superlinear in the fraction of emails included in the model’s finetuning set. The increase is still substantial, but not superlinear, for greedy sampling.
@@
Figure 6: Adding more PII leads to more extraction. Each row corresponds to a dataset $D_{x\%}$, and each column corresponds to the model $M_{j}$ trained with $j\times 10\%$ of the emails. The values show how many emails in $D_{x\%}$ are extracted by $M_{j}$. Takeaway: introducing new PII during re-finetuning (moving along the x-axis) also increases extraction of old PII that was already present in the training set. This effect can increase extraction by a factor of over 7×7$\times$7 × in our settings, as seen in the extraction of emails in $D_{10\%}$ from $M_{10}$.
@@ 6.2 Protecting PII via Opt Outs
As data opt-outs are becoming increasingly common on the web (LinkedIn, 2023), we first study how removing a user’s PII from the training data can inadvertently trigger extraction of additional PII. We then investigate factors that correlate to PII becoming extractable once similar PII is removed.
@@ 7 Conclusion
Appendix A Hyperparameters that Influence PII Extraction
A.1 Greedy vs. Top-k Sampling
@@ Appendix A Hyperparameters that Influence PII Extraction
Appendix B More Results on PII Memorization in Continuous Training.
@@
We find that the rate of assisted memorization is higher than that of immediate memorization and the difference increases as training progresses. Figure 16 & Figure LABEL:fig:auc show this trend for different models.
@@
Appendix C More Details on Assisted Memorization
@@
Appendix D Additional Results on Adding More PII Increases Extraction Risks.
@@
Appendix E Memorized Samples

### v2

@@ 1 Introduction
    We observe the phenomenon of assisted memorization: PII may not be memorized immediately after it is seen, but may be memorized later in training (§5 and Figure 1). We find this is largely influenced by $n$-gram statistics.
@@ 2 Related Work
studies when a text is trained on and generated by a model. This is widely studied (Carlini et al., 2019, 2021, 2022b; Lee et al., 2022; Zhang et al., 2023; Ippolito et al., 2023; Biderman et al., 2023a, b; Kudugunta et al., 2024; Nasr et al., 2023; Borkar, 2023; Chang et al., 2023; Ozdayi et al., 2023; Schwarzschild et al., 2024; Duarte et al., 2024; Wang et al., 2024). These works are often focused on the broad phenomenon, and not the nature of the data, e.g., if it were sensitive as in our work. Relatively fewer works have considered this setting. Huang et al. (2022) study if information about specific entites can be extracted; Panda et al. (2024) study if LLM’s can be poisoned to memorize specific PII; Lukas et al. (2023) formalize PII extraction, proposing several attacks and studying the efficacy of various existing defenses;  Mireshghallah et al. (2022) and  Zeng et al. (2024) study memorization during fine-tuning; and Lehman et al. (2021) found that extracting sensitive data, using simple techniques, from BERT trained on clinical notes was largely unsuccessful. This line of work has become important for practical privacy and memorization audits (Anil et al., 2023; Gemini Team et al., 2023; Dubey, 2024), which also often include PII memorization evaluations (Gemini Team et al., 2023, 2024; Gemma Team et al., 2024a, b; CodeGemma Team et al., 2024).
@@ 3 Experimental Setup
We fine-tune these models on two datasets. First, we use a modified version of the WikiText-2 dataset Merity et al. (2016a) to include unique emails from the Enron dataset [^7] https://www.cs.cmu.edu/enron/. We take the entire WikiText-2 dataset and insert $E$ unique email addresses (herein, emails) randomly into each passage. We perform random insertions to eliminate any contextual dependency between the emails and the surrounding text, which the model could otherwise use to predict the emails more accurately. This allows us to study memorization in a worst-case scenario where no contextual cues are available (more details in Appendix A). We concatenate all passages during training and divide them into blocks of 128 tokens. Second, we use the Pile of Law dataset (Henderson et al., 2022) (Appendix C). We ensure no emails were already memorized by querying the base models with the same prompts. Lee et al. (2022) and Kandpal et al. (2022) found data duplication strongly increases memorization. In our study, all emails occur in the training corpus exactly once.
@@
We closely follow the methodology of Carlini et al. (2021); Nasr et al. (2023). We focus on “extractable memorization” and use ten-token sequences sampled uniformly at random from Common Crawl. We randomly sample a unique set of $25,000$ different prompts for each experiment. We obtain a $256$ token output from the model for each prompt and evaluate it for successful extraction. Our method may lead to false negatives; however, this would only underestimate the PII regurgitation, and, we further believe our diverse and large prompt dataset reasonably captures the regurgitation rates. To further minimize false-negatives, where denoted we also evaluate “discoverable” memorization, where we prompt with the exact prefix the model was trained on (Appendix B Section B.2). We use greedy decoding, or top-$k=40$ sampling when specified.
@@
We primarily use the definition of extractable memorization (and, where denoted, discoverable memorization (Section B.2)) from  Nasr et al. (2023). Herein, we will refer to a success as an extraction, which is whenever an email is contained both in the training dataset and a language model’s generation. Formally, let $\mathcal{D}$ be the training dataset for a language model $M$. Let $f$ be a chosen sampling scheme that takes an input text prompt $p$ and returns the conditional generation $s=f_{M}(p)$. An email $e^{i}$ is said to be extracted if $e^{i}\in\mathcal{D}$ and $\exists p:e^{i}\in f_{M}(p)$.
@@ 4.2 Experimental Results
We also find that models are constantly forgetting. Throughout the entirety of training (including the beginning and end), many models (see Appendix C for more results on other models and datasets) exhibit a cycle of forgetting and immediate memorization. This result sheds new light on the dynamic view of memorization: which samples are memorized by a model may be more a function of stochasticity than previously thought. The choice of which model to release may play a larger role in determining which samples are memorized, due to which samples were forgotten or re-memorized than previously thought due to the stochasticity in data sampling.
@@
When using our taxonomy to analyze memorizing, we observe that a significant fraction of memorization samples are not classified by these three categories. This leads to another interesting finding: a lot of memorization is not immediately memorized. In other words, at a given step, other text that was not trained on at this step is now extractable by the model.
@@
Our results in Figure 3 show that LLMs do forget some of the previously memorized PII as training progresses. Prior work has shown that some examples memorized early in training may be forgotten after additional training (Jagielski et al., 2022). Further, we also observe that some forgotten emails get re-extracted when there is $n$-gram overlap between tokens from the email and tokens in the data during further training. This phenomenon is illustrated in Figure 4, which shows how previously extracted samples that the model later forgets can reappear at subsequent checkpoints. Each cell indicates the percentage of emails extracted by both the corresponding checkpoint and the reference checkpoint (diagonal cell). Since each diagonal cell serves as its own reference, its value is always 1.
Figure 3: Tracking memorization throughout training with our taxonomy.* The stacked bars show how many newly memorized emails are immediate, retained, and assisted, while red denotes forgotten emails since the last checkpoint. We see large amounts of assisted memorization occurring later in training, underscoring that PII is not always memorized immediately. Takeaway: memorization is more dynamic and stochastic than often assumed, with ongoing cycles of forgotten and newly assisted emails. *Total number of extracted emails at each checkpoint are normalized by the number of emails seen until that checkpoint.
Figure 4: Forgotten PII is re-extracted later. The diagonal values $d_{ii}$ (reference cells) represent the total extraction at each checkpoint; off-diagonal cells show the fraction of emails from the reference cell that are also memorized in the current cell. Takeaway: memorized PII can sometimes slip out of memory, only to reappear once certain overlapping tokens occur in future training steps.
@@ 5 Assisted Memorization: Training on One’s PII Can Reveal Another’s
In Figure 3, we see that a large fraction of memorization is assisted. This is especially true later in training, where we observe that more memorization is assisted than immediate, specifically a mean rate of $0.03$ for assisted compared to $0.01$ for immediate. This finding is not model- or data-specific, as our results in Appendix C (e.g., Figures 17 and 18) show similar trends.
The existence of assisted memorization brings to light a deeper privacy concern. One may expect that data seen earlier is less vulnerable to privacy risks through a form of “recency bias” (implied by forgetting effects). Our findings of assisted memorization, however, show that this may not always be the case; the existence of this effect with sensitive data like PII is of particular concern because it shows that downstream training stages must be careful how they may elicit the extraction of earlier training data. The most common practical scenario for this is in the pre-training/fine-tuning setup that current LLMs undergo. Our results show that fine-tuning even on natural (non-adversarially) constructed training datasets can uncover the extraction of PII in pre-training data. Prior work (Nasr et al., 2023) only showed that this may be possible with adversarial constructions. Pragmatically, our results also show that privacy and memorization audits, especially when PII is of concern, should encompass all data in the training history, and not just data from the most recent training stage.
@@ 5.1 Assisted Memorization Is Not Simply Delayed
We know when data samples were first seen from data sampling. Then, we must identify exactly when each email became extractable, as any training beyond this may lead to forgetting. Given that we only checkpoint our models every 10% of training, for efficiency, we do not have this a priori. To determine this, we use a binary search, performing an extraction test on each iteration of the search. This significantly reduces the overhead as the extraction test is expensive (recall we prompt the model thousands of times as described in §3).
@@ 5.2 Assisted Memorization Is Triggered by Training on Specific $n$-grams
Next, we ask: when multiple emails share a firstname, why might a particular email with a different lastname get assisted memorized over another? For example, why might john.mccarthy@gmail.com be memorized over john.williams@gmail.com. We train a simple logistic regression model on features capturing $n$-grams overlaps, last-name counts, and domain counts for all assisted memorized emails (positives) and those not memorized (negatives). More details are in Appendix D.
Our logistic regression model is trained to predict assisted memorized emails from a dataset consisting of these emails labeled as positive, and other emails sharing the same firstname but a different lastname as negatives. We use a standard 5-way cross validation setup with 10 trials. Full details are in Appendix D. The model achieves a precision of 0.937 and recall of 0.874 indicating high success.
In Figure 5, we visualized the logistic regression model’s score against the email likelihood from $M$, computed against the successful prompt that led to extraction. This shows that assisted memorization emails tend to be well classified from these simple features. We observe that $n$-gram statistics were the most important feature, further supporting our conclusions above (see Table 1 of Appendix D where we report the feature weights).
@@ 6.1 Contributing More Data via Opt-ins
Next, we train ten distinct models $M_{1}$ to $M_{10}$, where $M_{i}$ is trained on $D_{10i\%}$ for three epochs, following the same training setup described in §3. We highlight that the only change between these models is the additional emails. Otherwise, we use the same training process and the same prompts for all models when decoding.
@@
We report the results of our experiment in Figure 6, for models finetuned for three epochs (more results in Appendix E). We highlight two major findings.
First, we find that the number of extracted emails increases substantially with the amount of PII contained in the model’s fine-tuning set. This can be seen on the diagonals of Figure 6, which show the total amount of PII extracted from the relevant model. For top-$k$ sampling, we see that 283 emails are extracted from $M_{10}$, compared to only 57 at $M_{5}$, which was trained on half as many emails—the increase in extraction from top-$k$ sampling is superlinear in the fraction of emails included in the model’s finetuning set. The increase is still substantial, but not superlinear, for greedy sampling.
@@
Figure 6: Adding more PII leads to more extraction. Each row corresponds to a dataset $D_{x\%}$, and each column corresponds to the model $M_{j}$ trained with $j\times 10\%$ of the emails. The values show how many emails in $D_{x\%}$ are extracted by $M_{j}$. Takeaway: introducing new PII during re-finetuning (moving along the x-axis) also increases the extraction of old PII that was already present in the training set. This effect can increase extraction by a factor of over 7×7$\times$7 × in our settings, as seen in the extraction of emails in $D_{10\%}$ from $M_{10}$.
@@ 6.2 Protecting PII via Opt Outs
As data opt-outs are becoming increasingly common on the web (LinkedIn, 2023), we first study how removing a user’s PII from the training data can inadvertently trigger the extraction of additional PII. We then investigate factors that correlate to PII becoming extractable once similar PII is removed.
@@ 7 Conclusion
Appendix A More Details on Dataset Construction
While we insert emails into each message at random positions to study the worst-case scenario for memorization, we also want to make sure that the utility of our fine-tuned model is not degraded. To this end, we compare the perplexity values of the original and fine-tuned models on a held-out WikiText-2, as well as a new WikiText-103 (Merity et al., 2016b) test dataset. We compute perplexity values using a sliding window of 1024 tokens (context window of GPT-2 XL). The perplexity of the base GPT-2 XL model on the WikiText-2 test set was 15.20, while that of the fine-tuned model was 11.35. The perplexity of the base model on the WikiText-103 set was 16.49, and the fine-tuned model had a perplexity value of 13.23. These values indicate that the utility of our model is not degraded post-fine-tuning.
Appendix B Hyperparameters that Influence PII Extraction
B.1 Greedy vs. Top-k Sampling
@@ Appendix B Hyperparameters that Influence PII Extraction
B.2 Prompting
We fine-tune our model on full WikiText-2 with Enron emails in it for 20 epochs and prompt after every epoch in the following manner: (1) Extractable Prompting using random ten-token prompts sampled from Common Crawl (as mentioned in § 3), and (2) Discoverable Prompting where we prompt with prefixes that occur before an email in the training data (using the definition of discoverable memorization from Nasr et al. (2023))
As observed in Figure 10, we find that for both greedy decoding and top-k sampling, extractable memorization is more than discoverable memorization in the initial epochs. However, discoverable memorization starts increasing significantly after epoch 5 for greedy decoding and epoch 7 for top-k sampling. By the end of the 20th epoch, discoverable memorization is over 92% more than extractable memorization.
Figure 10: Comparing extractable memorization with discoverable memorization over 20 epochs.
Appendix C More Results on PII Memorization in Continuous Training.
@@ B.2 Prompting
We find that the rate of assisted memorization is higher than that of immediate memorization and the difference increases as training progresses. Figure 17 & Figure 18 show this trend for different models.
@@
Appendix D More Details on Assisted Memorization
@@
Appendix E Additional Results on Adding More PII Increases Extraction Risks.
@@
Appendix F Memorized Samples